TJ plans to erect 132 feet of fencing around his rectangular garden, using his house as one side of the garden. Each width (x) is 3 feet longer than twice the length (y). Find the dimensions of the garden.
Answer by checkley77(12569)

x=2y+3
2x+(2y+3)=132
2(2y+3)+2y+3=132
4y+6+2y+3=132
6y+9=132
6y=132-9
6y=123
y=123/6
y=20.5 ans. for the length.
x=2*20.5+3
x=41+3
x=44 ans for the width.
Proof:
2*44+(2*20.5+3)=132
88+41+3=132
132=132